# FLUID: Training-Free Face De-identification via Latent Identity Substitution

### v1

@@ Introduction
Linear editing is the simplest approach, where identity is modified by directly adding the editing direction to the original latent, as expressed in Equation 7. This formulation is common in latent editing literature (kwon2022diffusion; zhu2023boundary), and is especially effective for large semantic shifts, such as face de-identification.
@@
To maintain consistency with the latent data manifold, we implement a tangent-space traversal approach, inspired by studies on geodesic editing in generative latent spaces (hahm2024isometric; jeong2024training). Rather than applying direct linear addition, we project $\Delta h_{\text{id}}$ onto the tangent plane of the hypersphere at $z$. Let $z^{\text{unit}}=z/\|z\|$ and $\Delta h^{\text{unit}}_{\text{id}}=\Delta h_{\text{id}}/\|\Delta h_{\text{id}}\|$. The projection becomes:
$\Delta h_{\text{proj}}=\Delta h^{\text{unit}}_{\text{id}}-\langle\Delta h^{\text{unit}}_{\text{id}},z^{\text{unit}}\rangle\cdot z^{\text{unit}}.$ (12)
Then, the geodesically traversed latent is defined as:
$\hat{z}=\|z\|\cdot\left(\cos(\theta)\cdot z^{\text{unit}}+\sin(\theta)\cdot\Delta h_{\text{proj}}\right).$ (13)
Here, $0\leq\theta\leq\pi$ is an angle given by the norm of $h_{id}$ in radians. This way, we can still guide the latent walk’s direction and intensity, consider the geodesic manifold, and eliminate a hyperparameter (i.e., editing strength $\lambda$) at the same time. Linear edits offer simplicity and strong directionality, while tangent edits preserve manifold geometry for more localized and realistic transformations.
Figure 3: Qualitative comparison with SOTA face de-identification methods on in-domain (CelebA-HQ) and out-of-domain (FFHQ) images. FLUID produces de-identified face images with stricter attribute preservation than other methods. While CelebA-HQ aligns with the pretrained DM’s distribution, FFHQ highlights FLUID’s generalization ability to more diverse real-world inputs.
CelebA-HQ FFHQ
@@
CIAGAN 0.852 0.984 0.470 0.635 0.518 0.167 0.588 0.857 0.986 0.425 0.650 0.466 0.126 0.528
DiffAM 0.507 0.999 0.500 0.810 0.774 0.514 0.653 0.515 1.000 0.493 0.763 0.767 0.473 0.648
AMT‑GAN 0.492 0.999 0.653 0.855 0.787 0.658 0.681 0.408 1.000 0.668 0.841 0.821 0.638 0.623
DeepPrivacy2 0.696 1.000 0.536 0.897 0.458 0.530 0.714 0.652 1.000 0.519 0.802 0.366 0.479 0.662
Repaint 0.691 0.999 0.501 0.871 0.605 0.575 0.734 0.772 0.998 0.382 0.721 0.440 0.369 0.660
Diff-Privacy 0.921 0.982 0.543 0.686 0.586 0.407 0.756 0.901 0.987 0.504 0.711 0.536 0.320 0.711
Diffprivate 0.810 1.000 0.565 0.752 0.436 0.618 0.752 0.806 1.000 0.536 0.579 0.440 0.537 0.719
FLUID (tangent) 0.685 1.000 0.660 0.907 0.674 0.651 0.776 0.683 1.000 0.599 0.827 0.584 0.537 0.736
FLUID (linear) 0.739 1.000 0.620 0.883 0.646 0.638 0.786 0.697 1.000 0.581 0.799 0.562 0.512 0.730
Table 1: Quantitative performance comparison between FLUID and SOTA face de-identification methods across six key metrics. These include identity removal (SID), image quality (detect), and preservation of four facial attributes: emotion, gender, pose, and gaze. FLUID achieves improved balance between attribute preservation and de-identification, further enhancing realism beyond other SOTA methods.
Loss Functions
To guide this latent substitution process, we introduce a set of differentiable loss functions, which we term reagent losses. In analogy to SN1 reactions, where different chemical reagents interact with a reactive intermediate to influence the final product, each loss function acts as a distinct supervisory “reagent" that steers the optimization toward a desired transformation. Just as protic polar solvents in SN1 stabilize intermediates and facilitate leaving group departure, our identity loss promotes effective identity removal, while attribute and face mask losses stabilize the output by preserving semantic and spatial structure. Together, these loss functions form a tunable and composable framework that enables selective identity editing in the latent space without compromising image utility.
Identity Loss.
$\mathcal{L}_{\text{id}}=e^{-D_{\text{id}}(F_{x},F_{\hat{x}}))}$, (14)
Attribute Preservation Loss.
$\mathcal{L}_{\text{attr}}=D_{\text{KL}}(A_{x}||A_{\hat{x}})=\sum_{a=1}^{N_{\text{att}}}A_{x}^{(a)}\log\left(\frac{A_{x}^{(a)}}{A_{\hat{x}}^{(a)}}\right),$ (15)
Face Mask Loss.
To localize identity transformation and reduce artifacts, we apply a masked L1 loss using a facial region mask $M$ extracted via FaceXFormer’s face parser function. This helps restrict significant changes to face-related pixels, and can be written as:
$\mathcal{L}_{\text{mask}}=\sum_{i}|x_{i}-\hat{x}_{i}|(1-M_{i}),$ (16)
Total Objective.
These reagent losses are combined as:
$\mathcal{L}_{\text{total}}=\lambda_{id}\mathcal{L}_{\text{id}}+\lambda_{attr}\mathcal{L}_{\text{attr}}+\lambda_{mask}\mathcal{L}_{\text{mask}},$ (17)
where $\lambda_{id}=1.0$, $\lambda_{attr}=1.0$, and $\lambda_{mask}=0.5$ are weights that control the relative importance of each loss term. This weighted formulation enables fine-grained control over the trade-off between semantic fidelity and de-identification strength, offering a principled and flexible strategy for identity editing in a training-free diffusion framework.
Figure 4: Qualitative results of ablation study for loss function combinations. The top two rows display CelebA-HQ samples, while the bottom three show results on FFHQ.
Experiments
Baselines
To benchmark the effectiveness of our training-free facial identity editing approach, we compare it against SOTA facial de-identification models. Since no existing work addresses identity substitution in the h-space, we select representative baselines from two major categories: perturbation-based and semantic identity-shifting methods. Note that, unlike FLUID, several baselines rely on additional inputs such as reference faces, mask images, or extra training.
Perturbation-based methods.
We include high-fidelity perturbation models that apply visual transformations without semantic reconfiguration. Specifically, we evaluate against AMT-GAN (hu2022protecting) and DiffAM (sun2024diffam). These models, often based on makeup or region masking, are favored over traditional noise- and patch-based techniques for their superior visual realism.
Semantic identity-shifting methods.
We consider CIAGAN (maximov2020ciagan), DeepPrivacy2 (hukkelaas2023deepprivacy2), Repaint (lugmayr2022repaint), the edit variant of Diffprivate (le2025diffprivate), and Diff-Privacy (he2024diff), which directly synthesize alternative identities via generative priors. We use the official implementations and default hyperparameter settings provided in their respective publications for comparison. Our method uses a fixed learning rate of $lr=0.001$, an editing strength parameter of $\lambda=1000$, and a diffusion inversion timestep of $T=600$.
@@
w/ all 0.739 1.000 0.620 0.883 0.646 0.638
w/o $\mathcal{L}_{ID}$ 0.134 1.000 0.793 0.957 0.868 0.817
w/o $\mathcal{L}_{att}$ 0.827 0.999 0.549 0.793 0.590 0.616
w/o $\mathcal{L}_{mask}$ 0.820 0.999 0.653 0.902 0.555 0.591
Table 2: Quantitative results of ablation study based on CelebA-HQ edit results. Each loss function is omitted to assess the impact.
Datasets
Evaluation Metrics
Identity Deviation.
We define identity deviation as source identity distance (SID), which indicates the difference between the original and edited images in terms of identity. We use ArcFace (deng2019arcface), a SOTA angular margin-based FR model, to quantify SID. This is measured as the inverse cosine similarity between ArcFace embeddings of the original image $x$ and the de-identified result $\hat{x}$:
[…]
where $F_{Arc}(\cdot)$ denotes the ArcFace encoder.
Attribute Preservation.
To evaluate semantic consistency, we focus on gender, emotion, pose, and gaze. Gender and emotion are predicted using DeepFace (serengil2021hyperextended), and their preservation metric for each attribute $a\in{\text{gender},\text{emotion}}$ is:
[…]
where $\mathbb{I}[\cdot]$ is the indicator function.
To evaluate pose consistency, we use the pose estimator function from FaceXFormer and compute a deviation-based preservation score that combines pitch, yaw, and roll angular differences into a single metric. Given pose deviations $\Delta_{\text{pitch}}$, $\Delta_{\text{yaw}}$, and $\Delta_{\text{roll}}$ between original and de-identified images, we first calculate the root mean square (RMS):
$\text{RMS}_{\text{pose}}=\sqrt{\frac{\Delta_{\text{pitch}}^{2}+\Delta_{\text{yaw}}^{2}+\Delta_{\text{roll}}^{2}}{3}}.$ (20)
The pose preservation score is then normalized as:
$R_{\text{pose}}=\max\left(0,1-\frac{\text{RMS}_{\text{pose}}}{\tau}\right),$ (21)
where $\tau=5$ represents the angle threshold in degrees for significant pose deviation. Higher scores indicate better pose preservation, with perfect preservation yielding $R_{\text{pose}}=1$. Inspired by Disguise (cai2024disguise), we include gaze as an additional utility using L2CS-Net (abdelrahman2023l2cs), estimating pitch and yaw. The gaze preservation score reuses Equations 20 and 21, with $\tau=15$.
Detectability.
To assess visual plausibility and face detectability, we count the number of face detection successes $R_{\text{det}}$. We use both Dlib (king2009dlib) and MTCNN (zhang2016joint) and average the counts.
Overall Metric.
We use a nested harmonic mean to prevent an aggregate from over‑valuing a configuration that inflates one attribute while degrading another. First, we form a harmonic mean score of four attributes:
$\mathrm{HM}_{\text{attr}}=\frac{4}{\frac{1}{R_{\text{emotion}}}+\frac{1}{R_{\text{gender}}}+\frac{1}{R_{\text{pose}}}+\frac{1}{R_{\text{gaze}}}},$ (22)
then apply another harmonic mean with three elements, $R_{\text{id}}$, $R_{\text{det}}$, and $\mathrm{HM}_{\text{attr}}$ to obtain the overall score $\mathrm{HM}_{\text{overall}}$. Because harmonic mean is dominated by low components, a method cannot mask weak privacy or a degraded attribute with a single strong metric; nesting enforces internal balance among attributes before combining the three pillars (privacy, detectability, semantics) concisely.
Results
Quantitative and qualitative comparison.
Table 1 and Figure 3 present FLUID’s quantitative and qualitative performance against existing methods. Notably, despite requiring no task-specific training and minimal attribute supervision, FLUID achieves the best overall balance between identity suppression and attribute preservation.
On CelebA-HQ, tangent edit yields the highest gender preservation (0.907), alongside strong emotion (0.660) and gaze (0.651) consistency, resulting in a high balance score of 0.776. While linear edits show higher SID (0.739), they introduce off-manifold artifacts, which reduce visual realism and semantic stability. In contrast, tangent-space editing, by respecting the local geometry of the h-space, enhances generalization to out-of-distribution data, as evidenced in FFHQ where it achieves the highest overall score (0.736). This underscores the strength of our training-free framework, matching or exceeding domain-specific methods without retraining.
Identity disentanglement validation.
Tangent vs. linear editing.
Why does tangent editing outperform linear editing, and does it generalize? In high-dimensional latent spaces, linear edits often drift off the data manifold, resulting in degraded outputs. Tangent editing, conversely, constrains the edit direction to the local tangent plane of $h$, preserving norm and curvature (as shown in Equations 12 and 13). This geodesic formulation avoids semantic drift, significantly improving attribute preservation, visual quality, and cross-domain generalization, particularly in challenging settings like emotion and pose changes (Table 1).
Figure 5: Identity-attribute trade-off curve across increasing linear edit strengths.
Ablation Study
To evaluate the individual contribution of the proposed reagent loss components, we conduct an ablation study, summarized in Table 2 and Figure 4. Removing $\mathcal{L}_{\text{ID}}$ severely compromises identity suppression, underscoring its critical role in effective de-identification. Excluding $\mathcal{L}_{\text{att}}$ leads to significant semantic drift, particularly altering key attributes such as gender and facial expression, thereby violating attribute preservation. When $\mathcal{L}_{\text{mask}}$ is omitted, we observe increased background artifacts and reduced visual realism, despite slight improvements in some numerical scores, highlighting a trade-off between perceived quality and structural integrity.
Limitations and Societal Impact
While our method excels in identity suppression, attribute preservation, and visual quality, it has some limitations. Its iterative latent-space optimization hinders real-time use in resource-limited settings. Performance may drop on highly out-of-distribution data, given its design assumptions and reliance on CelebA-HQ-style priors. Also it may not reliably preserve high-frequency details (e.g., text, glasses, occlusions). First, as a training-free method relying on latent-space optimization, it requires iterative backpropagation, which may hinder real-time deployment in resource-constrained environments. Second, although our approach generalizes across domains, its effectiveness may decline on highly out-of-distribution data, given its design assumptions and reliance on CelebA-HQ-style priors. Third, the model may not consistently preserve high-frequency details (e.g., text, glasses, occlusions), which are not explicitly optimized in our loss formulation.
Conclusion
Supplementary Material
In this supplementary material, we provide the following:
  • Detailed explanation on SN1, the substitution reaction in organic chemistry that we used as a metaphor.
  • Algorithm flow of our method.
  • Experiments and source of hyperparameter selection.
  • User study on human identification of de-identified faces.
  • Failure cases, especially produced by FFHQ results.
The Metaphor
In organic chemistry, a substitution reaction refers to the process of replacing one functional group in a molecule with another, while largely preserving the rest of the molecular structure (rizzo_sn1sn2; ashenhurst_sn1sn2). Among such reactions, the SN1 (unimolecular nucleophilic substitution) mechanism proceeds in two distinct stages. First, the leaving group spontaneously departs, resulting in the formation of a highly reactive, electron-deficient carbocation intermediate. This unstable intermediate reflects a vulnerable yet reorganizable state. In the second step, a nucleophile selectively attacks the carbocation, completing the substitution and stabilizing the molecule.
Figure 6: The SN1 mechanism illustrated in (lowry2012sn1).
In contrast, the SN2 (bimolecular nucleophilic substitution) mechanism is a single-step reaction where the nucleophile directly replaces the leaving group at a fixed site, without forming an intermediate. SN2 is deterministic and rigid, typically operating only when the substitution site is clearly accessible and unambiguous.
This distinction forms the conceptual foundation of our approach. We draw a parallel between molecular substitution and identity editing in diffusion models. We conceptualize the latent representation of a facial image as a molecular structure in which identity acts as a substitutable functional group, embedded within a stable semantic backbone encoding attributes such as pose, expression, lighting, and age. Just as substitution reactions allow targeted replacement of molecular components while preserving overall structure, we aim to selectively replace identity in latent space while retaining all non-identity semantics.
Our method aligns naturally with the SN1 mechanism. First, we do not assume the identity subspace is predefined. Instead, we dynamically discover and suppress identity directions in the latent space via loss-guided optimization, analogous to the formation of the carbocation intermediate. Once identity is removed, we get an anonymized identity vector, akin to a nucleophile completing the substitution. This two-step, context-aware, and optimization-driven process mirrors the SN1 mechanism in both structure and intent. Moreover, unlike SN2, our method does not rely on a fixed, interpretable site for substitution, but instead adapts to the distributional geometry of the latent space, reinforcing the relevance of the SN1 analogy. This metaphor offers a conceptually intuitive and technically grounded framework for controlled, semantic-preserving identity substitution in the latent space of diffusion models.
Procedure of FLUID
On top of Figure 2 in main paper, we provide our method’s flow using pseudocodes in Algorithm 1. Our method starts with initialization, where an identity embedding, an attribute distribution, and a face mask are obtained by auxiliary models given an input image. A direction vector $\Delta h_{\text{id}}$ is initialized randomly as a vector with a small norm $m$. Then $\Delta h_{\text{id}}$ and the original image’s vector $z$ are fused by either linear (Equation 7) or tangent edit (Equations 12 and 13), producing an edited vector $\hat{z}$. The optimization steps proceed by repeatedly calculating losses between the input images and the edited image (Equation 17), made by decoding $\hat{z}$ (Equation 8), while backpropagating the loss values to $\Delta h_{\text{id}}$ to correct it (Equation 11). After a given number of steps $n_{opt}$, the final vector $\hat{z}$ is decoded through the DM’s denoising process, yielding a de-identified image $\hat{x}$. For all random initializations and possible operations, our random seed is fixed with the value of $1006$.
Inspired by spherical interpolation’s success in other generative models, we therefore traverse edits along local tangent planes, controlling the angular step size to stay within the observed shell. As Figure 7 and Table 1 show, this angular traversal yields fewer artifacts and better quality than straight linear edits, which can wander significantly beyond the typical norm range into areas where the decoder produces unrealistic outputs.
Figure 7: Comparison of linear against tangent edits in the h-space. We fit a 3-dimensional principal component analysis (PCA) pool with $5000$ vectors sampled from $100$ editing trajectories, and project the desired vectors onto this PCA space to visualize their trajectory. (a) Linear edits apply raw $\Delta h_{\text{id}}$, allowing the vector norm to drift beyond the typical shell and producing visual artifacts. (b) Tangent edits project $\Delta h_{\text{id}}$ onto the local tangent plane at a constant norm, constraining the trajectory and reducing artifacts.
Hyperparameter Selection and Analysis
Reagent loss weights.
We conducted ablation studies on loss weights, and the results are reported in Table 4. The cases of 0 for each weight are excluded because the results are presented in Table 2. $\lambda_{\text{id}}=0.5$ and $\lambda_{\text{id}}=1.5$ each lead to the weakest and strongest SID scores, making the identity loss trivial and dominant, respectively. Lowering $\lambda_{\text{attr}}$ results in lower attribute preservation rates but higher SID due to weaker constraints. However, both damage the balances, which are proven by low overall scores. Increasing $\lambda_{\text{mask}}$ forces the background information to be fixed firmly and limits the editable region for de-identification. In this case, the optimization process decides to alter the gender information to compensate for the difficulty. Although $\lambda_{\text{mask}}=0.5$ shows a slightly lower balance score than $\lambda_{\text{id}}=1.5$, we select it for superior emotion preservation at the cost of moderately lower SID. This configuration ($\lambda_{\text{mask}}=0.5$) is applied to the reported experiments in the main paper.
Denoising steps and computational analysis.
Similarly, we conduct another ablation study to select the number of denoising steps. As written in Table 3, increasing the number of reverse steps tend to reduce SID and retain attributes, especially gaze. Although 8 denoising steps show the best balance with the strongest SID, we select 16 steps for our experiments for improved stability and consistency, ensuring reliable performance across diverse inputs and more attribute preservation. Even though the number of denoising steps is much less than that of DDPM, thanks to the implementation of deterministic denoising by DDIM (song2020denoising), naively backpropagating through the iterative reverse steps still requires considerable amount of memory. Therefore, we implement gradient checkpointing (chen2016training) to significantly reduce memory consumption, enabling us to run the implementation on our hardware, a single Titan XP GPU with 12GB of memory. With the help of gradient checkpointing, FLUID runs with approximately 6GB of memory and takes under 3 seconds for each optimization step, including all auxiliary model operations.
$p_{\theta}^{(t)}(x_{t-1}\mid x_{t})=\mathcal{N}\left(\sqrt{\alpha_{t-1}}\,\mathbf{P}_{t}\left(\epsilon_{\theta}^{t}(x_{t})\right)+\mathbf{D}_{t},\sigma_{t}^{2}\mathbf{I}\right).$ (23)
Linear edit strength.
In the main paper, we clarify that the editing strength $\lambda$ for linear edits is set to $1000$. In contrast, Boundary Diffusion offers the editing strength within the range of [150,150][-150,150], as its direction vector encodes the signed distance to a hyperplane (i.e., boundary) defined by a linear SVM. Since we do not rely on such boundaries during optimization, and the norm of $\Delta h_{\text{id}}$ varies throughout the process, we conduct an ablation study on the linear editing strength $\lambda$ in the range of $100$ to $1000$, as shown in Table 5. Unlike other ablation studies where trade-offs are evident, we observe trivial trade-off patterns in this case. Larger values of $\lambda$ consistently lead to increased SID, with constantly high detectability scores and slightly decreasing attribute preservation, thereby improving the overall score. We attribute the increase in SID to a larger exploration range enabled by stronger scaling factors within a limited number of optimization steps. Based on this observation, we select $\lambda=1000$ as our default linear edit strength.
Input: Input image $x$, pretrained decoder $D$, weights $\lambda_{\text{id}},\lambda_{\text{attr}},\lambda_{\text{mask}}$, step size $\eta$, editing type (linear or tangent), initialized $\Delta h_{\text{id}}$’s norm $m$, number of optimization steps $n_{\text{opt}}$.
@@ Conclusion
 $\Delta h_{\text{id}}\leftarrow\Delta h_{\text{id}}-\eta\cdot\nabla_{\Delta h_{\text{id}}}\mathcal{L}_{\text{total}}$ $\triangleright$ Equation 11

### v2

@@ 4.2. Latent Navigation: From Euclidean to Geodesic
Linear editing is the simplest approach, where identity information is modified by directly adding the editing direction to the original latent, as expressed in Equation 7. This formulation is widely used in latent editing literature (kwon2022diffusion; zhu2023boundary), and is particularly effective for inducing large semantic shifts, such as face de-identification.
@@
We observe that the manifold of the h-space forms a thin hypershell that closely resembles a hypershpere. Although the h-space appears to support linear edibility, we hypothesize that this behavior arises because subsequent normalization operations force linearly edited vectors back to the manifold. However, linear edits that momentarily escape the manifold without any consideration of its intrinsic geometry, introduce undesirable artifacts that persist even after normalization. Simply normalizing the linear edited latent to match the original norm (i.e., $z||$) remains unreliable, as this procedure still inherits the distortions cause by ‘cutting through the manifold’. Therefore, to maintain consistency and to appreciate the latent data manifold, we implement a tangent-space traversal strategy, inspired by recent work on geodesic editing in generative latent spaces (hahm2024isometric; jeong2024training). Rather than applying direct linear addition, we first project identity-editing direction $\Delta h_{\text{id}}$ onto the tangent plane of the hypersphere at $z$ to enable edits remain on the surface of the manifold. Let $z^{\text{unit}}=z/\|z\|$ and $\Delta h^{\text{unit}}_{\text{id}}=\Delta h_{\text{id}}/\|\Delta h_{\text{id}}\|$. The projection becomes:
(9) $\Delta h_{\text{proj}}=\Delta h^{\text{unit}}_{\text{id}}-\langle\Delta h^{\text{unit}}_{\text{id}},z^{\text{unit}}\rangle\cdot z^{\text{unit}}.$
Then, the geodesically traversed latent is defined as an automated spherical interpolation (slerp). To automatically perform slerp, we start with the basic form:
(10) $\text{slerp}(p_{0},p_{1};\Omega)=\frac{\text{sin}((1-\Omega)\cdot\Psi)}{\text{sin}(\Psi)}\cdot p_{0}+\frac{\text{sin}(\Omega\cdot\Psi)}{\text{sin}(\Psi)}\cdot p_{1},$
where $p_{0}$ and $p_{1}$ are the two points on the arc, $\Psi$ is the angle between $p_{0}$ and $p_{1}$, and $\Omega$ is the degree of interpolation between $p_{0}$ and $p_{1}$. Since the two vectors $z^{\text{unit}}$ and $\Delta h_{\text{proj}}$ are perpendicular to each other, $\Psi=90$. Therefore,
(11) $\text{slerp}(z^{\text{unit}},\Delta h_{\text{proj}};\Omega)$ $=\frac{\text{sin}((1-\Omega)\cdot 90\degree)}{\text{sin}(90\degree)}\cdot z^{\text{unit}}+\frac{\text{sin}(\Omega\cdot 90\degree)}{\text{sin}(90\degree)}\cdot\Delta h_{\text{proj}}$
@@
Finally, with $\theta=\Omega\cdot 90\degree$, we have our final optimization objective as:
(12) $\hat{z}=\|z\|\cdot\left(\cos(\theta)\cdot z^{\text{unit}}+\sin(\theta)\cdot\Delta h_{\text{proj}}\right).$
Here, we substitute $\theta$ with an angle given by the norm of $h_{id}$ in radians. This way, we can still guide the latent walk’s direction and intensity, consider the geodesic manifold, and eliminate a hyperparameter (i.e., editing strength $\lambda$) at the same time. Linear edits offer simplicity, while tangent edits preserve manifold geometry for more localized and realistic transformations.
Input: Input image $x$, pretrained decoder $D$, weights $\lambda_{\text{id}},\lambda_{\text{attr}},\lambda_{\text{mask}}$, step size $\zeta$, editing type (linear or tangent), initialized $\Delta h_{\text{id}}$’s norm $m$, number of optimization steps $n_{\text{opt}}$.
@@
 $\Delta h_{\text{id}}\leftarrow\Delta h_{\text{id}}-\zeta\cdot\nabla_{\Delta h_{\text{id}}}\mathcal{L}_{\text{total}}$
@@
4.3. Loss Functions
To guide this latent substitution process, we introduce a set of differentiable loss functions that steer the optimization toward a desired transformation. These loss functions form a tunable and composable framework that enables selective identity editing in the latent space while preserving the overall utility and semantic consistency of the reconstructed image.
Identity Loss.
(13) $\mathcal{L}_{\text{id}}=e^{-D_{\text{id}}(F_{x},F_{\hat{x}}))}$,
Attribute Preservation Loss.
(14) $\mathcal{L}_{\text{att}}=D_{\text{KL}}(A_{x}||A_{\hat{x}})=\sum_{a=1}^{N_{\text{att}}}A_{x}^{(a)}\log\left(\frac{A_{x}^{(a)}}{A_{\hat{x}}^{(a)}}\right),$
Face Mask Loss.
To localize identity transformation and reduce artifacts, we apply a masked L1 loss using a facial region mask $M$ extracted via FaceXFormer’s face parser function. Note that we omit the final thresholding step that discretizes each pixel’s probability to 0 or 1, and obtain the raw probabilities of pixels which belong to facial components. This helps restrict significant changes to pixels related to face components (e.g., eyebrows, nose, lips) and smoothens the boundary splitting facial components and non-facial components, reducing boundary-like artifacts seen in inpainting-based de-identification approaches (lugmayr2022repaint; hukkelaas2019deepprivacy; hukkelaas2023deepprivacy2). The loss function can be written as:
(15) $\mathcal{L}_{\text{mask}}=\sum_{i}|x_{i}-\hat{x}_{i}|(1-M_{i}),$
Total Objective.
The three losses are combined as:
(16) $\mathcal{L}_{\text{total}}=\lambda_{id}\mathcal{L}_{\text{id}}+\lambda_{att}\mathcal{L}_{\text{att}}+\lambda_{mask}\mathcal{L}_{\text{mask}},$
where $\lambda_{id}=1.0$, $\lambda_{att}=1.0$, and $\lambda_{mask}=0.5$ are weights that control the relative importance of each loss term. This weighted formulation enables fine-grained control over the trade-off between semantic fidelity and de-identification strength, offering a principled and flexible strategy for identity editing in a training-free diffusion framework.
4.4. Procedure of FLUID
Combining the components introduced in the methodology, we summarize FLUID’s overall workflow using pseudocode in Algorithm 1. Our method starts with initialization, where an identity embedding, an attribute distribution, and a face mask are obtained by auxiliary models given an input image. A direction vector $\Delta h_{\text{id}}$ is initialized randomly as a vector with a small norm $m$. Then $\Delta h_{\text{id}}$ and the original image’s vector $z$ are fused by either linear edit (Equation 7) or a tangent edit (Equations 9 and 12), producing an edited latent vector $\hat{z}$. The optimization steps proceed by repeatedly calculating losses between the input images and the edited image (Equation 16), made by decoding $\hat{z}$ (Equation 8). These loss values are backpropagated to update $\Delta h_{\text{id}}$, progressively refining the identity-editing direction. After a designated number of steps $n_{opt}$, the final latent vector $\hat{z}$ is decoded through the DM’s denoising process, yielding a de-identified image $\hat{x}$. Inspired by spherical interpolation’s success in other generative models, we therefore traverse edits along local tangent planes, controlling the angular step size to stay within the observed shell. As Figure 8 and Table 1 show, this angular traversal yields fewer artifacts and less utility damage than straight linear edits, which may drift beyond the typical norm range into regions where the decoder produces unrealistic or unstable outputs.
5. Experiments
5.1. Baselines
To benchmark the effectiveness of our training-free facial identity editing approach, we compare it against SOTA semantic shift-based face de-identification models. We report FALCO (barattin2023attribute)  [^1] We report FALCO on CelebA-HQ following the authors’ official evaluation protocol (CelebA-HQ split and attribute-based evaluation). Extending FALCO to other datasets requires dataset-specific engineering and, in some cases, additional dataset information or annotations to enable comparable attribute-based metrics. As such extensions may introduce protocol-dependent variations and raise concerns about inconsistent evaluation, we omit multi-dataset FALCO results. This choice is consistent with prior works (kung2025face; kung2025nullface; he2024diff; yang2024g), which also report FALCO primarily on CelebA-HQ., DeepPrivacy2 (hukkelaas2023deepprivacy2), Repaint (lugmayr2022repaint), the edit variant of Diffprivate (le2025diffprivate), and FAMS (kung2025face), which directly synthesize alternative identities via generative priors. We use the official implementations and default hyperparameter settings provided in their respective publications for comparison. Our method uses a fixed learning rate of $lr=0.0005$, an editing strength parameter of $\lambda=1000$, and a diffusion inversion timestep of $T=600$ with Adam optimizer (kingma2014adam).
5.2. Datasets
5.3. Evaluation Metrics
Identity Deviation.
We define identity deviation as source identity distance (SID), which measures the difference between the original and edited images in terms of identity. To quantify SID while measuring the transferability of de-identification across various FR models, we compute this metric using three popular FR models; ArcFace (deng2019arcface), SphereFace (liu2017sphereface), and MobileFaceNet (chen2018mobilefacenets). SID is measured as the inverse cosine similarity between a pair of face embeddings in the FR feature space, each belonging to the original image $x$ and the de-identified result $\hat{x}$: